OpenStudy (anonymous):

divide. Show your work. 3n^2-n/n^2-1 divide n^2/n+1

Nnesha (nnesha):

\[\huge\rm \frac{ 3n^2 -n }{ n^2 -1 } \div \frac{ n^2 }{ n+1 }\] first you should factor 3n^2 -n and n^2 -1

Nnesha (nnesha):

when you change division to multiplication you should flip 2nd fraction

OpenStudy (amorfide):

just stop helping this guy, he has posted about 20 questions which require the same method but won't attempt them himself after being shown how to do it over and over

OpenStudy (bibby):

what's with the toxic attitude kid

Nnesha (nnesha):

\[\huge\rm \frac{ a }{ b } \div \frac{ c }{ d } = \frac{ a }{ b} \times \frac{ d }{ c }\] for example like this

Nnesha (nnesha):

thanks :)

OpenStudy (anonymous):

wc

Nnesha (nnesha):

wc ?

OpenStudy (anonymous):

hahaha I mean welcome

Nnesha (nnesha):

okay o^_^o

Nnesha (nnesha):

got it ??

OpenStudy (anonymous):

yeah I have the answer thanks